OpenStudy (anonymous):

Gerhard deposited $5,600 into a savings account that earns 4.5% simple interest each year calculated annually. What is the future value of Gerhard's account after 12 years? A. $14,224 B. $8,624 C. $3,024 D. $5,852

OpenStudy (mathstudent55):

I = Prt I = interest earned P = present value r = annual interest rate as decimal t = time in years The future value is F = P + I or F = P + Prt

OpenStudy (stamp):

\[i=$5,600(4\%/year)12\ years\]=$3,024\[5600+3024=\]